3. Suppose XYZ Company is a dominant firm in a particular industry. The demand curve for this industry’s product is ? = 200 − 10?, where Q is the quantity demanded and P is the price. The supply curve for the small firms in the industry is ?? = 20 + 2?, where ?? is the total amount supplied by all the small firms combined. XYZ Company’s marginal cost is ?? = 2??, where ?? is XYZ Company’s output.

Question: Calculate the price of the product produced by XYZ Company to maximize its profit.
Answer #1

Qs=20+2P

MC=2QA. Firm produces at a point where P=MC=2QA

Qa=P/2.

Total quantity supplied=Qs+Qs=20+5/2P

and total Qd=200-10P

Equilibrium is when Qd=Qs

200-10P=20+5/2P

180=12.5P

price=180/12.5=14.4 and firm produces when P=MC

2Qa=14.4

Thus QA=7.2 and price=14.4 in order to maximise profit